# TimeLDM: Latent Diffusion Model for Unconditional Time Series Generation

### v1

@@ Abstract
Time series generation is a crucial research topic in the area of deep learning, which can be used for data augmentation, imputing missing values, and forecasting. Currently, latent diffusion models are ascending to the forefront of generative modeling for many important data representations. Being the most pivotal in the computer vision domain, latent diffusion models have also recently attracted interest in other communities, including NLP, Speech, and Geometric Space. In this work, we propose TimeLDM, a novel latent diffusion model for high-quality time series generation. TimeLDM is composed of a variational autoencoder that encodes time series into an informative and smoothed latent content and a latent diffusion model operating in the latent space to generate latent information. We evaluate the ability of our method to generate synthetic time series with simulated and realistic datasets, benchmark the performance against existing state-of-the-art methods. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series. Sores from Context-FID and Discriminative indicate that TimeLDM consistently and significantly outperforms current state-of-the-art benchmarks with an average improvement of 3.4$\times$ and 3.8$\times$, respectively. Further studies demonstrate that our method presents better performance on different lengths of time series data generation. To the best of our knowledge, this is the first study to explore the potential of the latent diffusion model for unconditional time series generation and establish a new baseline for synthetic time series.
[…]
123
1 Introduction
In recent years, time series generation has served a pivotal role in a wide range of applications, such as forecasting, imputing missing values, data augmentation, and simulation in unseen scenarios. Moreover, generating time series data can be a valuable approach to solving the complex challenges associated with data privacy concerns. It allows organizations to unlock the value of their data assets while maintaining the highest standards of privacy protection and regulatory compliance. Numerous studies have used various architectures of deep neural networks for synthetic realistic time series data, including Variational Autoencoder (VAE) based methods [19, 15], Generative Adversarial Network (GAN) based methods [13, 14, 16, 17], Flow-based methods [18] and more [28, 29]. Typically, GAN-based methods have gained plenty of attention from researchers. For instance, Cot-GAN [17] incorporates a specialized loss function based on a regularized Sinkhorn distance, which originates from the principles of causal optimal transport theory. TimeGAN [13] merges the adaptability of the unsupervised GAN framework with the control afforded by supervised training in autoregressive models. However, despite the potential application, the generation results are still unsatisfactory from simulated and realistic time series data.
Figure 1: Distribution of our TimeLDM and Diffusion-TS on multivariate datasets. The blue solid line is for original data, and the yellow dotted line is for synthetic data. TimeLDM aligns better with the original data.
@@ 1 Introduction
Most recently, diffusion models (DMs) [1, 2] have shown impressive performance on image generation [41, 40] and beyond [39, 38]. DMs construct a diffusion process that gradually degrades the training data structure, and learned neural networks to reverse this contaminated data by progressive denoising. To further improve the synthesis quality of time series data, several studies have also applied DMs for time series synthesis, such as DiffTime [3] adopts future mixup and autoregressive initialization as a condition to generate time information. Diffusion-TS [4] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models. Those methods have emerged as a superior learning architecture to GANs in generative modeling. However, existing approaches often apply diffusion models directly in the data space, which typically consists of limited information and noisy features. Therefore, we are interested in searching for a more flexible framework for time series modeling.
@@
Figure 2: Structure of our proposed TimeLDM. (a) shows the components of TimeLDM, consisting of transformer encoder, reparameterization, diffusion process, reverse process, and transformer decoder. (b) shows the detail of transformer encoder and reparameterization. (c) shows the architecture of latent diffusion model. (d) shows the detail of transformer decoder.
@@
Latent space generation, an efficient alternative model in generative architectures, adopts a pre-trained autoencoder to transfer the generation tasks from the input space to a greater flexible latent domain. In this paper, inspired by the success of the diffusion model on latent space [21, 23], we propose an efficiently synthesize time series method to overcome the above limitations by adopting a smoother and informative latent presentation, named TimeLDM (Time Latent Diffusion Model). As shown in Figure 2 (a), We first transform the raw time series data into an embedding space and train the encoder and decoder network for the VAE. The well-studied VAE converts the time series data into the smoothed and informative latent space. After training the VAE, we apply the encoder to extract the latent vectors from the training set and apply them as the target of the latent diffusion model, which is designed with a denoising MLP. During inference, we generate the latent vectors from the latent diffusion model and then apply the VAE decoder to synthesize the time series.
@@
We validate the performance of our proposed approach for different benchmarks, including simulated and realistic time series datasets. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series. As shown in Table 2, the scores from Context-FID and Discriminative indicate that TimeLDM consistently and significantly outperforms current state-of-the-art benchmarks. For example, TimeLDM achieves an average improvement of 3.8$\times$ in the Discriminative score. Furthermore, we demonstrate that our TimeLDM presents better performance on different lengths of time series data generation, Table 4 presents the results.
The main contributions of this paper are summarized as follows:
@@
  • We propose TimeLDM, a latent diffusion-based method that leverages the high-fidelity image synthesis ability into unconditional time series generation. To the best of our knowledge, this is the first work to explore the potential of LDM for unconditional time series generation and establish a new baseline for synthetic time series.
@@
  • We evaluate the ability of our method to synthetic time series with simulated and realistic datasets, benchmark the performance against existing state-of-the-art methods. Empirically, TimeLDM shows better performance than existing time series generation methods both qualitatively and quantitatively. The ablation study presents the proposed loss function of VAE which plays a crucial role in improving the capability of TimeLDM.
@@
  • Furthermore, we evaluate our method with different lengths of time series data, TimeLDM presents better performance on the proposed benchmark datasets compared with current state-of-the-art methods.
2 Related Works
2.1 Time Series Generation
@@
In a broad range of fields, diverse deep generative models have demonstrated their ability to generate high-quality samples, where generating time series stands as a particularly challenging endeavor within limited information and noisy features. Early methods based on Generative Adversarial Networks [12] (GANs) have been extensively investigated for time series generation. The GAN comprises two neural networks that are trained concurrently: a discriminator trained to differentiate between real samples from the training dataset and those that are generated, and a generator that is trained to produce samples that “fool” the discriminator. For example, TimeGAN [13] applies an embedding function and supervised loss to the original GAN for capturing the temporal dynamics of data throughout time. RTSGAN [14] and PSA-GAN [16] generate long univariate time series samples of high quality with self-attention mechanism. Cot-GAN [17] incorporates a specialized loss function based on a regularized Sinkhorn distance, which originates from the principles of causal optimal transport theory. While Variational Autoencoders [20] (VAEs) and Flows [26, 27] , also drawn attention on researchers. For instance, Fourier Flows [18] is proposed as a method based on normalizing flows followed by a chain of spectral filters leading to an exact likelihood optimization. TimeVAE [19] implements an interpretable temporal structure and achieves reasonable results on time series synthesis.
Recent research [5, 3, 4] has been exploring the use of diffusion models [1] to generate time series, developing on the successes of forward and reverse processing in other areas such as images [8], video [9], text [10], and audio [11]. Among them, TimeGrad [5] is a conditional diffusion model that predicts in an autoregressive manner, with the denoising process guided by the hidden state of a recurrent neural network. CSDI [6] uses self-supervised masking to guide the denoising process like image inpainting. SSSD [7] based on the structured state space model. DiffTime [3] approximated the diffusion function based on CSDI where they remove the side information provided as embedding. They also introduced guided DiffTime to handle new constraints such as trend or fixed-value, without re-training. Diffusion-TS [4] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models.
2.2 Generative Modeling in the Latent Space
@@
Although generative models in the data space have achieved significant success, the latest emerging latent generative models have demonstrated several advantages, including more compact and disentangled representations, robustness to noise, and greater flexibility in controlling generated styles. To enable faster training and inference on limited computational resources. Due to the compaction and smoothness of the latent space, latent generative models are preferred for their computational efficiency. For instance, the recent models [21, 23] have achieved great success in image generation as they exhibit better scaling properties and expressivity than the vanilla diffusion models in the data space. The success of the latent diffusion model in image generation has also inspired their applications in video [24], audio [25], tabular [35], and text [37] domains. In this paper, we explore the application of the latent diffusion models for unconditional time series generation tasks.
3 Method
In this section, we formally introduce TimeLDM, a latent diffusion model (LDM) for unconditional time series generation. Figure 2 presents the overall structure of our TimeLDM, which consists of an encoder-decoder module for variational autoencoder (VAE), reparameterization trick for latent information sampling, and the LDM. In Section 3.1, we formulate the time series generation task first. Then, we introduce the details of VAE in Section 3.2. After that, we unveil the LDM from network architecture to mathematical formulation in Section 3.3. Finally, we summarize the training and sampling procedures in Section 3.4.
3.1 Problem Statement
3.2 Time Series Autoencoding
To overcome the weakness of data domain generation, we are interested in presenting the time series signals $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$ into an informative and smoothed latent space. The latent representation is $\mathbf{L}=\left\{\mathbf{z}\right\}_{i=1}^{N}$, where $\mathbf{z}_{1:\tau}=\left(z_{1},\ldots,z_{\tau}\right)\in\mathbb{R}^{\tau\times m}$ denotes the latent feature and $m$ is the dimension of representation. The framework of VAE is shown in Figure 2 (b) and (d). As we can see, it designs with encoder and decoder module, VAE $=(\mathcal{E}_{\phi}(\mathbf{x}),\mathcal{D}_{\xi}(\mathbf{z}))$, where the encoder $\mathcal{E}_{\phi}$ learn the latent variable $\mathbf{z}=\mathcal{E}_{\phi}(\mathbf{x})$ and the decoder $\mathcal{D}_{\xi}$ decode latent feature $\mathbf{z}$ back to data domain $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$. Unlike the vanilla VAE apply the ELBO loss function to train, here we adopt $\beta$-VAE [32], where the coefficient $\beta$ adaptively balances the reconstruction loss and KL-divergence loss for effective training.
@@ 3.2 Time Series Autoencoding
         Get embedding pattern
         Get positional pattern
@@
VAE’s Encoder. As shown in Figure 2 (b), the VAE’s Encoder $\mathcal{E}_{\phi}$ first apply one dimension convolutional neural network to learn an embedding pattern $\mathbf{e}=emb(\mathbf{x})\in\mathbb{R}^{\tau\times m}$ from the temporal structures $\mathbf{x}_{1:\tau}\in\mathbb{R}^{\tau\times d}$, then a learnable positional encoding $pe\in\mathbb{R}^{\tau\times m}$ equip to the embedding feature for adaptively learning the time series positional information. After that, we train two transformer encoders to learn the mean $\mu\in\mathbb{R}^{\tau\times m}$ and log variance $\sigma\in\mathbb{R}^{\tau\times m}$ from the positional encoding feature $\mathbf{e}^{pe}_{1:\tau}=\mathbf{e}_{1:\tau}+pe$, respectively. Next, we obtain the latent variables $\mathbf{z}_{1:\tau}$ from the reparameterization trick Function 1.
@@
VAE’s Decoder. As shown in Figure 2 (d), the aim of VAE’s Decoder $\mathcal{D}_{\xi}$ is to minimize the reconstruction error by generating outputs that closely resemble the original time series information. The input to the Decoder $\mathcal{D}_{\xi}$ consists of latent variables sampled from a typically Gaussian distribution, which is derived from the Encoder using the reparameterization trick. The architect of VAE decoder is designed with a transformer decoder that incorporates both self-attention and cross-attention mechanisms. The input also respects the learning embedding and positional encoding process. Finally, it generates the realistic samples of time series data $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$.
Training Loss. The training objective of the VAE consists of the reconstruction loss $\mathcal{L}_{\text{recon}}$ and the KL divergence $\mathcal{L}_{\text{KL}}$. The reconstruction loss is composed of the $\mathcal{L}_{1}$ norm and $\mathcal{L}_{2}$ norm in the data domain, as well as the $\mathcal{L}_{1}$ norm in the frequency domain [42] between the input data and the reconstructed signal. Inspire by [29] accurate reconstruction time series signals with Fourier-based loss term, we denote Fast Fourier Transformation loss term $\|\mathcal{FF}\mathcal{T}(\mathbf{x}),\mathcal{FF}\mathcal{T}(\hat{\mathbf{x}})\|$ to learn frequency domain information of time series. $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weights to balance three losses.
@@
KL divergence loss regularizes the mean and log variance of the latent space. As shown below 4, the $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ is probabilistic output from the encoder $\mathcal{E}_{\phi}$ that represents the approximate posterior of latent variable $\mathbf{z}$ given the input $\mathbf{x}$; $\mathcal{N}(\mathbf{z};\mu,\sigma)$ is the prior on $\mathbf{z}$ The $\beta$ is adaptively tuned during training, where $\beta=\lambda\beta,\lambda<1$. If the $\mathcal{L}_{\text{recon}}$ fails to decrease with defined steps, the $\beta$ will decrease to encourage the model to pay more attention to the reconstruction term $\mathcal{L}_{\text{recon}}$.
@@
Finally, the overall training objective of the VAE is as below. For the adaptive $\beta$, we set $\beta_{\max}=0.01$, $\beta_{\min}=10^{-5}$, and $\lambda=0.7$, where the $\beta_{\max}$ is initial setting, and $\beta_{\min}$ is the minimum number of the adaptive $\beta$.
@@
3.3 Latent Diffusion Model
After the preparations mentioned above, a trained VAE allows us to access an informative and smoother latent space $\mathbf{L}=\{\mathbf{z}\}_{i=1}^{N}$. Consequently, we are ready to design a diffusion model to generate the latent space information. Figure 2(c) presents the neural network architecture of LDM. First, we reshape the sampling representation into one dimension before passing through a linear layer. Next, we transform the time step $t$ into sinusoidal embeddings $\boldsymbol{t}_{\text{emb}}$, and added to the $\text{Linear}(\mathbf{z})$ After that, we apply four linear layers to learn the denoising pattern. Finally, we reshape the latent representation back to the input shape. Following [2], we adopt below forward process 5 and reverse process 6 to obtain noising data and learn to reverse back:
@@ 3.3 Latent Diffusion Model
3.4 Training and Sampling

### v2

@@ Abstract
Time series generation is a crucial research topic in the area of decision-making systems, which can be particularly important in domains like autonomous driving, healthcare, and, notably, robotics. Recent approaches focus on learning in the data space to model time series information. However, the data space often contains limited observations and noisy features. In this paper, we propose TimeLDM, a novel latent diffusion model for high-quality time series generation. TimeLDM is composed of a variational autoencoder that encodes time series into an informative and smoothed latent content and a latent diffusion model operating in the latent space to generate latent information. We evaluate the ability of our method to generate synthetic time series with simulated and real-world datasets and benchmark the performance against existing state-of-the-art methods. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series. For example, TimeLDM achieves new state-of-the-art results on the simulated benchmarks and an average improvement of 55% in Discriminative score with all benchmarks. Further studies demonstrate that our method yields more robust outcomes across various lengths of time series data generation. Especially, for the Context-FID score and Discriminative score, TimeLDM realizes significant improvements of 80% and 50%, respectively. The code will be released after publication.
I INTRODUCTION
Time series generation holds a pivotal role across numerous applications, such as robotics [1, 2], autonomous driving [3, 4], and healthcare [5, 6]. Additionally, generating time series can be a valuable approach to solving the complex challenges associated with data privacy concerns. It enables agents to learn a wealth of information without containing any actual sensitive data, providing a safer framework for model training and development.
@@ I INTRODUCTION
Numerous studies have used various architectures of deep neural networks for synthetic realistic time series data, including Variational Autoencoder (VAE) based methods [7, 8], Generative Adversarial Network (GAN) based methods [9, 10, 11, 12], and Diffusion-based methods [13, 14]. Typically, Diffusion-based methods have gained plenty of attention from researchers. For instance, DiffTime [13] adopts future mix-up and autoregressive initialization as a condition to generate time information. Diffusion-TS [14] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models. Those methods have emerged as a superior learning architecture in generative modeling to others. However, existing approaches often apply learning models directly in the data space, which typically consists of limited information and noisy features. Therefore, we are interested in searching for a more flexible framework for time series modeling.
Figure 1: t-SNE visualization on the stocks dataset, TimeLDM shows better overlap between the generated data and original data than TimeVAE.
@@
Figure 2: Structure of our proposed TimeLDM. (a) shows the components of TimeLDM, consisting of the transformer encoder, reparameterization, diffusion process, reverse process, and transformer decoder. (b) shows the details of the transformer encoder and reparameterization. (c) shows the architecture of the latent diffusion model. (d) shows the details of the transformer decoder.
@@
Latent space generation, an efficient alternative model in generative architectures, adopts a pre-trained autoencoder to transfer the generation tasks from the input space to a greater flexible latent domain. In this paper, inspired by the success of the diffusion model on latent space [15, 16], we propose an efficiently synthesized time series method to overcome the above limitations by adopting a smoother and informative latent presentation, named TimeLDM (Time Latent Diffusion Model). As shown in Figure 2 (a), We first transform the raw time series data into an embedding space and train the encoder and decoder network for the VAE. The well-studied VAE converts the time series data into the latent space. After that, we apply the latent information as the target of the latent diffusion model (LDM), which is designed with a denoising MLP. During inference, we generate the latent vectors from the LDM and then apply the VAE decoder to synthesize the time series.
@@
We validate the performance of our proposed approach for different benchmarks, including simulated and real-world time series datasets. Qualitatively and quantitatively, we find that the proposed TimeLDM persistently delivers high-quality generated time series (see Figure~5). In Table II and III, the Discriminative scores of TimeLDM consistently outperform current state-of-the-art benchmarks. Furthermore, Table IV demonstrates that TimeLDM presents better performance on different lengths of time series data generation. The main contributions of this paper are summarized as follows:
@@
  • We propose TimeLDM, a latent diffusion-based method that leverages the high-fidelity image synthesis ability into unconditional time series generation. To the best of our knowledge, this is the first work to explore the potential of LDM for unconditional time series generation.
@@
  • We evaluate the ability of our method to synthetic time series with simulated and real-world datasets. Empirically, TimeLDM shows better performance than existing generation methods both qualitatively and quantitatively. The ablation study presents the proposed loss function of VAE, which plays a crucial role in improving the capability of our method.
@@
  • Furthermore, we evaluate TimeLDM with different lengths of time series data, which presents better performance on the proposed benchmark datasets compared with current state-of-the-art methods.
II RELATED WORKS
@@
Time Series Generation. Deep generative models have demonstrated their ability to generate high-quality samples across a wide array of fields, where generating time series stands as a particularly challenging endeavor within limited information and noisy features. Early methods based on GANs [17] have been extensively investigated for time series generation. For example, TimeGAN [9] applies an embedding function and supervised loss to the original GAN for capturing the temporal dynamics of data throughout time. Cot-GAN [12] incorporates a specialized loss function based on a regularized Sinkhorn distance, which originates from the principles of causal optimal transport theory. While VAEs [18] also drew the attention of researchers. For instance, TimeVAE [7] implements an interpretable temporal structure and achieves reasonable results on time series synthesis. Recent research [13, 14] has been exploring the use of diffusion models [19] to generate time series, developing on the successes of forward and reverse processing in other areas such as images [20], video [21], text [22], and audio [23]. Among them, DiffTime [13] approximated the diffusion function based on CSDI [24] where they remove the side information provided as embedding. Diffusion-TS [14] combines the interpretability component, such as trend and multiple seasonality, to model time series using denoising diffusion models.
@@
Generative Modeling in the Latent Space. Although generative models in the data space have achieved significant success, the latest emerging LDMs [15, 16] have demonstrated several advantages, including more compact and disentangled representations, robustness to noise, and greater flexibility in controlling generated styles. LDMs have achieved great success in image generation as they exhibit better scaling properties and expressivity than the vanilla diffusion models in the data space. The success of the LDM in image generation has also inspired their applications in video [25], audio [26], tabular [27], and text [28] domains. In this paper, we explore the application of the LDM for unconditional time series generation tasks.
@@
         Get the embedding pattern
         Get the positional pattern
@@
III METHOD
TimeLDM, as shown in Figure 2, consists of an encoder-decoder module for VAE, a reparameterization trick for latent information sampling, and the LDM. In this section, we formulate the time series generation task first. Then, we introduce the details of VAE and LDM from network architecture to mathematical formulation. Finally, we summarize the training and sampling procedures.
III-A Problem Statement
III-B Time Series Autoencoding
To overcome the weakness of data domain generation, we are focusing on presenting the time series signals $\mathbf{S}=\left\{\mathbf{x}\right\}_{i=1}^{N}$ into an informative and smoothed latent space. The latent representation is $\mathbf{L}=\left\{\mathbf{z}\right\}_{i=1}^{N}$, where $\mathbf{z}_{1:\tau}=\left(z_{1},\ldots,z_{\tau}\right)\in\mathbb{R}^{\tau\times m}$ denotes the latent feature and $m$ is the dimension of representation. The framework of VAE is shown in Figure 2 (b) and (d). As we can see, it designs with encoder and decoder module, VAE $=(\mathcal{E}_{\phi}(\mathbf{x}),\mathcal{D}_{\xi}(\mathbf{z}))$, where the encoder $\mathcal{E}_{\phi}$ learn the latent variable $\mathbf{z}=\mathcal{E}_{\phi}(\mathbf{x})$ and the decoder $\mathcal{D}_{\xi}$ decode latent feature $\mathbf{z}$ back to data domain $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$. Here we adopt $\beta$-VAE [29], the coefficient $\beta$ adaptively balances the reconstruction loss and KL-divergence loss for effective training.
VAE’s Encoder. As shown in Figure 2 (b), the VAE’s Encoder $\mathcal{E}_{\phi}$ first apply a convolutional neural network to learn an embedding pattern $\mathbf{e}=emb(\mathbf{x})\in\mathbb{R}^{\tau\times m}$ from the temporal structures $\mathbf{x}_{1:\tau}\in\mathbb{R}^{\tau\times d}$, then a learnable positional encoding $pe\in\mathbb{R}^{\tau\times m}$ equip to the embedding feature for adaptively learning the time series positional information. After that, we train two transformer encoders to learn the mean $\mu\in\mathbb{R}^{\tau\times m}$ and log variance $\sigma\in\mathbb{R}^{\tau\times m}$ from the positional encoding feature $\mathbf{e}^{pe}_{1:\tau}=\mathbf{e}_{1:\tau}+pe$, respectively. Next, we obtain the latent variables $\mathbf{z}_{1:\tau}$ from the reparameterization trick Function 1.
@@ III-B Time Series Autoencoding
VAE’s Decoder. As shown in Figure 2 (d), the aim of VAE’s Decoder $\mathcal{D}_{\xi}$ is to minimize the reconstruction error by generating outputs that closely resemble the original time series information. The input to the Decoder $\mathcal{D}_{\xi}$ consists of latent variables sampled from a typically Gaussian distribution, which is derived from the Encoder using the reparameterization trick. The architect of the VAE decoder incorporates both self-attention and cross-attention mechanisms. The input also respects the learning embedding and positional encoding process. Finally, it generates the realistic samples of time series data $\hat{\mathbf{x}}=\mathcal{D}_{\xi}({\mathbf{z}})$.
Training Loss. The training objective of the VAE consists of the reconstruction loss $\mathcal{L}_{\text{recon}}$ and the KL divergence $\mathcal{L}_{\text{KL}}$. The reconstruction loss is composed of the $\mathcal{L}_{1}$ norm, $\mathcal{L}_{2}$ norm in the data domain, and the Fast Fourier Transformation (FFT) [30] loss term $\|\mathcal{FF}\mathcal{T}(\mathbf{x}),\mathcal{FF}\mathcal{T}(\hat{\mathbf{x}})\|$ in the frequency domain [31], which is inspired by HyperTime [32] for accurate time series reconstruction. $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weights to balance three losses.
@@
KL divergence loss regularizes the mean and log variance of the latent space. As shown Equation 3, the $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ is probabilistic output from the encoder $\mathcal{E}_{\phi}$ that represents the approximate posterior of latent variable $\mathbf{z}$ given the input $\mathbf{x}$; $\mathcal{N}(\mathbf{z};\mu,\sigma)$ is the prior on $\mathbf{z}$ The $\beta$ is adaptively tuned during training, where $\beta=\lambda\beta,\lambda<1$. If the $\mathcal{L}_{\text{recon}}$ fails to decrease with defined steps, the $\beta$ will decrease to encourage the model to pay more attention to the reconstruction term.
@@
Finally, the overall training objective of the VAE is as below. For the adaptive $\beta$, we set $\beta_{\max}=10^{-2}$, $\beta_{\min}=10^{-5}$, and $\lambda=0.7$, where the $\beta_{\max}$ is initial setting, and $\beta_{\min}$ is the minimum number of the adaptive $\beta$.
@@
III-C Latent Diffusion Model
After the preparations mentioned above, a trained VAE allows us to access the latent space $\mathbf{L}=\{\mathbf{z}\}_{i=1}^{N}$. Figure 2(c) presents the neural network architecture of LDM. First, we reshape the sampling representation into one dimension before passing through a linear layer. Next, we transform the time step $t$ into sinusoidal embeddings $\boldsymbol{t}_{\text{emb}}$, and added to the $\text{Linear}(\mathbf{z})$ After that, we apply four linear layers to learn the denoising pattern. Finally, we reshape the latent representation back to the input shape. Following [33], we adopt below forward process Equation 5 and reverse process Equation 6 to obtain noising data and learn to reverse back:
Parameter Sines MuJoCo Stocks ETTh fMRI
dim(x) 5 14 6 7 50
Attention heads 2 2 2 2 2
Attention head dimension 16 16 16 16 16
Encoder layers 1 1 2 2 1
Decoder layers 2 2 3 3 2
Batch size 1024 1024 512 1024 1024
Hidden dimension of LDM 1024 4096 1024 1024 4096
TABLE I: Hyperparameters of VAE and LDM.
III-D Training and Sampling